OpenStudy (anonymous):

Find Real Solutions to the following by GRAPHING. 1.) -15x^3 + 5X^2 + 6X 2.) 9 + X^2 = 6X

OpenStudy (aum):

Click on the graph to enlarge it: http://awesomescreenshot.com/0f93nrbt12

OpenStudy (anonymous):

Ok great! what about 9 + X^2 = 6X

OpenStudy (aum):

For 1) the real roots are: -0.487, 0, 0.821 2. 9 + x^2 = 6x x^2 - 6x + 9 = 0 Go to www.desmos.com and type in x^2-6x+9. It will graph it. The points where it cuts or touches the x-axis are the real solutions.

OpenStudy (anonymous):

Your awesome man thanks again!

OpenStudy (aum):

You are welcome.